OpenStudy (anonymous):

would (3.6X10^5) x (7.2x10^7) come out to 25.92X10^12?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

\[(3.6*10{}^{\wedge}5) (7.2*10{}^{\wedge}7)=2.592*10{}^{\wedge}13 \]

OpenStudy (anonymous):

rob , why would the decimal move?

OpenStudy (anonymous):

gah. i dont even know why im second guessing your answer!

OpenStudy (anonymous):

@savannahw88 it depends entirely on what canonical form you've chosen to represent the number. Since contextually it appears to be a scientific notation problem, you always want to keep a single whole part in the one's place.

OpenStudy (anonymous):

\[25.92*10{}^{\wedge}12=2.592 * 10{}^{\wedge}13 \]

OpenStudy (anonymous):

yes. scientific notation. sorry i wasnt specific.

OpenStudy (anonymous):

well here it goes. I had five questions and just wanted to seek a second opinion on this last one. thanks guys. or gals.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

wait one more double check. (-3p)^2 simplified is -3p^2.....

OpenStudy (anonymous):

\[(-3p){}^{\wedge}2\text{ //}\text{Simplify} \to 9 p^2 \]

OpenStudy (anonymous):

the same would go for (5p)^-2?

OpenStudy (anonymous):

25p^2?

OpenStudy (anonymous):

\[(5p){}^{\wedge}-2=\frac{1}{(5p){}^{\wedge}2}=\frac{1}{25 p^2} \]

OpenStudy (anonymous):

is that because the exponent is negative?

OpenStudy (anonymous):

Yes.